To maximize profit, the firm produces units of output per day, charges a price of $, and earns a total profit ofFill in the blanks using the information in the following graph of a monopoly: MC ATC AVC Demand MR 200 250 285 Output per da
Answer #1

Since , it is given that the firm is a monopoly, we know that the profit maximizing condition would be the output where MC= MR,

At this point , the firm produces 200 units of output and the corresponding price is $45. The cost corresponding to this output i.e. ATC corresponding is $32.

Hence the profit in this case will be : (P- ATC)Q

= (45-32)(200) = 13(200) = 2600